# Kilonova Light Curve Parameter Estimation Using Likelihood-Free Inference

### v2

@@ Abstract
We present a parameter estimation algorithm on kilonova light curves using likelihood-free inference. Our inference is optimized through a pre-trained embedding network that marginalizes the time of arrival and the luminosity distance of the signal. We find that parameter inference utilizing a pre-trained embedding outperforms the use of likelihood-free inference alone, reducing training time and offering the capability to marginalize over certain nuisance parameters. The model is capable of retrieving the intrinsic parameters of the kilonova light curves with a comparable accuracy and precision to nested sampling methods while taking significantly less computational time. This framework has been integrated into the publicly available Nuclear Multi-Messenger Astronomy codebase so users can leverage the model for their inference purposes. This algorithm is broadly applicable to parameterized or simulated light curves of other transient objects.
@@
pubyear: 2024pagerange: Kilonova Light Curve Parameter Estimation Using Likelihood-Free InferenceB
@@ 1 Introduction
For several decades, the mergers of compact binaries made up of two neutron stars (BNS), or a neutron star with a black hole (NSBH), have been theorized to emit a sufficient amount of neutron star ejecta to supply heavy elements through a process known as rapid nucleosynthesis, commonly referred to as the r-process (Lattimer & Schramm, 1974; Eichler et al., 1989; Freiburghaus et al., 1999). With the observation of the well-studied kilonova AT2017gfo (Tanvir et al., 2017) and the associated gravitational wave GW170817 and gamma ray burst GRB170817A (Abbott et al., 2017), recent research has been done on estimating the properties of the neutron star ejecta (Kasen et al., 2017; Bulla, 2023). This has led to developments in understanding the properties and equation of state of neutron stars (Radice et al., 2018; Mathias et al., 2024), additional measurements of the Hubble constant (Bulla et al., 2022), and investigating the contributions of these mergers to heavy element production via the r-process (Siegel, 2019).
@@
Over the past few years, developments in machine learning (ML) techniques have explored various avenues in reducing the computational cost and time for studying light curves. While ML-based improvements have been made in the vitesse of light curve generation using variational autoencoders (Saha et al., 2024) and neural networks (Peng et al., 2024), parameter estimation of this data remains an untapped field. One disadvantage of sampling methods is that the likelihood calculation of $p(d|\theta)$ at each inference step is computationally expensive. To avoid this calculation, likelihood-free or simulation-based inference (LFI/SBI) offers a potential alternative method (see Cranmer et al. (2020) for a review).
In this paper, we describe our framework for faster parameter inference using LFI with comparable accuracy and precision to nested sampling (Skilling, 2006). We simulate light curves using the Ka2017 model grid and ZTF noise realizations. We use an embedding neural network for the light curves to marginalize over the distance to the source and the exact peak time over a data segment. This allows us to focus on the intrinsic parameters of the light curve such as $M_{ej}$, $V_{ej}$, and $X_{lan}$, which are further described in Section 2. This is achieved by creating alternate, or shifted, views of each light curve with the same intrinsic parameters, while varying the time and luminosity distance. These views are jointly embedded by minimizing the variance, invariance, and covariance of the two data views. Conditioned on this data summary, a normalizing flow is used to infer these kilonova parameters on a logarithmic scale: $\log_{10}(M_{ej}/M_{\odot})$, $\log_{10}(V_{ej}/c)$, and $\log_{10}(X_{lan})$. We compare our posterior distributions to nested sampling results produced with the PyMultiNest package (Feroz et al., 2009). We explain our selection of model and priors for the data in Section 2. Section 3 details the ML architecture and demonstrates its efficacy on simulated kilonova light curves, with the results provided in Section 4.
@@ 2 Data
In this section, we discuss prior selection and data preparation for the ML model. We simulate light curve data using the Ka2017 hydrodynamical grids, which are available through the NMMA package (Kasen et al., 2017; Pang et al., 2024). We focus on obtaining realistic light curves viewable from ZTF, thereby simulating data in all three of its bands: ztfg, ztfr, and ztfi.
The composition and appearance of a kilonova is controlled by the spectral energy distribution (SED). The Ka2017 model parametrizes this SED through the combined dynamical and disk mass ($M_{ej}$), velocity ($V_{ej}$), and lanthanide fraction ($X_{lan}$) of the ejecta. Lanthanide fraction, $X_{lan}$, can serve as a metric for heavy r-process matter, as lanthanides contribute anywhere from 1 to 10 % of the total ejecta mass (Kasen et al., 2017). The line transitions from lanthanides greatly increases the opacity, resulting in long diffusion times and longer light curves. The Ka2017 model accounts for two kilonova mass ejection methods: a brief stripping of the stars due to tidal forces ($10^{-3}M_{\odot}<M_{ej}<10^{-2}M_{\odot}$), and a secondary ejection of mass from the accretion disk blown by neutrino-driven winds ($10^{-2}M_{\odot}<M_{ej}<10^{-1}M_{\odot}$) (Martin et al., 2015). The initial shock occurs at high speeds on the order of a dynamical velocity $V_{dyn}\sim 0.2c-0.3c$, while the disk mass disperses at rates of $V_{disk}\sim 0.05c-0.1c$. These quantities also affect the peak of the light curve, with large ejecta mass and velocity producing brighter and longer lasting kilonova signals. Ka2017 combines these quantities as the ejecta parameters $M_{ej}$and $V_{ej}$, which we use to parametrize our model. In this study, we neglect the effects of host galaxy and Milky Way extinction effects on the light curve.
@@
To ensure that light curves are visible by ZTF, we selected broad parameter ranges that were capable of producing detections in addition to the considerations provided by Kasen et al. (2017). These priors are detailed in Table 1. We choose time and luminosity distance, $(t,d_{L})$, as our nuisance parameters that we wish to marginalize. Our time prior is selected to create a conservative variability for where the light curve peaks. Our distance prior allows our model to be sensitive to the edge of EM detectablility with ZTF and the expected two detector LIGO O4 BNS inspiral range of 25-130 Mpc (Abbott et al., 2018). While GW170817 was detected at 40 Mpc, optical surveys are expected to be sensitive to kilonovae at a distance of up to 200 Mpc with a limiting magnitude of 22 mag (Sagués Carracedo et al., 2021).
@@
Our data is simulated at a regular cadence of one detection every six hours over a 20 day period of time. While this is an optimistic detection scheme, we discuss potential relaxation schemes in Section 4. We sample uniformly from our intrinsic parameters $\log_{10}(M_{ej}/M_{\odot})$, $\log_{10}(V_{ej}/c)$, and $\log_{10}(X_{lan})$, and our extrinsic parameters $t$ and $d_{L}$. In the case where the light curve is dim or short-lived, non-detections are encoded as the magnitude detection limit of 22.0 in all bands. This is a typical conservative restriction given that ZTF is limited to 20.8, 20.6, and 19.9 mag [21.6, 21.4, and 21.9 mag] for the g, r, and i bands for a 30s [300s] exposure time (Bellm et al., 2019; Almualla et al., 2021). Since the broad priors produce both extremely short and long light curves, we pad the data with our non-detection limit in order to maintain a standard length of data for our model. For training purposes, a minimum amount of 8 cumulative detections in any of the three bands is set as a requirement.
@@
To train the network, we create a set of “fixed" data, $d$, which share the same initial arrival time $t=0$ and a luminosity distance of $d_{L}=50$ Mpc. For each fixed light curve, an alternative “shifted" light curve $d^{\prime}$ is generated by modifying $t$ and $d_{L}$ in accordance with the priors in Table 1. A single fixed light curve will have its shifted counterpart that shares the precise values for $\log_{10}(M_{ej}/M_{\odot})$, $\log_{10}(V_{ej}/c)$, and $\log_{10}(X_{lan})$, resulting in the same physical phenomena but a different observed light curve. In this way, two views of data generated by the same intrinsic parameters are created. Additionally, each light curve is simulated with unique instrument noise to ensure that all views of the same intrinsic parameters are distinct. We utilize both views to train our data summary, and the shifted view to train our normalizing flow.
@@ 3 Method
In this section, we present our ML-based framework for parameter estimation using LFI. We introduce the concept of an embedding network that permits the marginalization of certain parameters and reduces the size of the following inference network in Section 3.1. We then use normalizing flows to perform parameter inference, which is described in Section 3.2.
@@ 3.1 Data Summary
In order to perform LFI using normalizing flows, we choose parameters ${\theta}_{i}\sim p(\mn@boldsymbol{\theta})$, and simulate light curves $\mathbf{\mathbf{d}}_{i}$. The pairs ${\theta}_{i},\mathbf{\mathbf{d}}_{i}\}$ are used to learn the posterior directly by maximizing the log-likelihood,
$\log\mathcal{L}=\sum_{i}p(\mn@boldsymbol{\theta}_{i}|\mathbf{\mathbf{d}}_{i}).$ (1)
However, a key challenge in implementing LFI on high dimensional data such as light curves is an effective data summary. In this case, we would like to perform inference on the intrinsic parameters, like $M_{ej}$, $V_{ej}$, or $X_{lan}$, even if the data depends on other extrinsic parameters. In the case of a kilonova light curve, signals can be seen as unique if they arrive at different times or occur at different distances, even when the intrinsic parameters of the ejecta are the same. An LFI network unaware of these translations requires the extrinsic parameters of time and distance to be inferred alongside the other physical parameters or marginalized over, leading to a larger, more expensive network. To mitigate the size of the network, we can marginalize over the nuisance parameters by creating an appropriate data summary that preserves the information of only the intrinsic parameters, which are of interest in this case. We utilize self-supervised learning (SSL) to accomplish this goal.
This technique has been used previously in Chatterjee et al. (2023) for simpler, lower dimensional signal morphologies, demonstrating that the creation of an embedding space aids in faster parameter inference. The application of self-supervised learning to create a summary extends to compressing cosmological data and encoding symmetries of partial differential equations as seen in Akhmetzhanova et al. (2024); Mialon et al. (2023). Additionally, self-supervised learning has been used to simplify the complexity of gravitational waves as an initial denoising step (Lin & Kong, 2024). Rodrigues & Gramfort (2020) demonstrates a different approach (YuleNet) for data summarization for LFI with timeseries data. As a result, we also implement a summarization step for our high-dimensional light curve data.
@@
Our full network architecture is explained in Figure 2’s flowchart. After the representation has been extracted, we pass $\gamma$ and $\gamma^{\prime}$ to our 20-dimensional expander layer, $h$, such that the outputs $x,x^{\prime}$ are produced:
@@
Here, $x$ and $x^{\prime}$ represent the data in an embedding space. We call this the data summary. To correlate $x$ and $x^{\prime}$, we utilize the VICReg loss function (Bardes et al., 2021), which is calculated as a weighted sum:
@@
$\begin{split}L_{VICReg}(x,x^{\prime})&=\lambda_{1}MSE(x,x^{\prime})+\lambda_{2}[\sqrt{Var(x)+\epsilon}\\ &+\sqrt{Var(x^{\prime})+\epsilon}]+\lambda_{3}[C(x)+C(x^{\prime})],\end{split}$ (4)
@@
where we choose equal weights $\lambda_{1}=\lambda_{2}=\lambda_{3}=1$ for this study. We minimize the VICReg similarity loss between the embeddings $x$ and $x^{\prime}$ such that the regularized variance and covariance of the batches are individually reduced, while the invariance between batches is reduced through mean-squared error.
@@
To properly encode the effects of changing $t$ and $d_{L}$, we must create two views of the data. We repetitively create 50 light curves using one identical combination of our physical parameters ($\log_{10}(M_{ej}/M_{\odot})$, $\log_{10}(V_{ej}/c)$, and $\log_{10}(X_{lan})$) with unique noise instances. This is done iteratively with novel combinations of the physical parameters to produce a dataset. Each of light curves in this fixed set are positioned at a $d_{L}$ of 50 Mpc, and start at the same time. Our second data view is created by adjusting $t$ and $d_{L}$ of each light curve of the fixed set according to the priors listed in Table 1. Our training dataset contains 8,729 unique combinations of $M_{ej}$, $V_{ej}$, and $X_{lan}$, resulting in 872,900 total light curves. The data representation, $\gamma$, is trained jointly using these sets of fixed and shifted light curves. 10 minutes of processing time is required to train the network for 50 epochs on one NVidia V100 GPU.
Figure 3: In this figure, we see four combinations of $\log_{10}(M_{ej}/M_{\odot})$, $\log_{10}(V_{ej}/c)$, and $\log_{10}(X_{lan})$plotted in orange, green, red, and purple. Each set contains 1000 light curves that have been shifted in time and distance. We observe that irrespective of their time and distance, the light curves are clustered with those that share the same $\log_{10}(M_{ej}/M_{\odot})$, $\log_{10}(V_{ej}/c)$, and $\log_{10}(X_{lan})$. This indicates that our nuisance parameters ($t,d_{L}$) have been marginalized.
@@
Figure 3 displays the representation $\gamma^{\prime}$ for four combinations of physical injection parameters to showcase various areas of the prior selection. We observe that after training, light curves of different $\log_{10}(M_{ej}/M_{\odot})$, $\log_{10}(V_{ej}/c)$, and $\log_{10}(X_{lan})$are clustered together even with different values of $t$ and $d_{L}$, indicating that the two invariant parameters have been successfully marginalized. We utilize the data summarization network as a step in our normalizing flows training.

### v3

# Rapid Parameter Estimation for Kilonovae Using Likelihood-Free Inference

@@ Abstract
Rapid parameter estimation is critical when dealing with short lived signals such as kilonovae. We present a parameter estimation algorithm that combines likelihood-free inference with a pre-trained embedding network, optimized to efficiently process kilonova light curves. Our method is capable of retrieving the mass, velocity, and lanthanide fraction of the neutron star ejecta with an accuracy and precision on par with nested sampling methods while taking significantly less computational time. Our inference uniquely utilizes a pre-trained embedding network that marginalizes the time of arrival and the luminosity distance of the signal, allowing inference of signals at distances up to 200 Mpc. We find that including a pre-trained embedding outperforms the use of likelihood-free inference alone, reducing training time, model size, and offering the capability to marginalize over certain nuisance parameters. This framework has been integrated into the publicly available Nuclear Multi-Messenger Astronomy codebase, enabling the broader scientific community to deploy the model for their inference purposes. Our algorithm is broadly applicable to parameterized or simulated light curves of other transient objects, and can be adapted for quick sky localization.
@@
pubyear: 2025pagerange: Rapid Parameter Estimation for Kilonovae Using Likelihood-Free Inference9
@@ 1 Introduction
For several decades, the mergers of compact binaries made up of two neutron stars (BNS), or a neutron star with a black hole (NSBH), have been theorized to emit a sufficient amount of neutron star ejecta to supply heavy elements through the rapid neutron-capture process, commonly referred to as the r-process (Lattimer & Schramm, 1974; Eichler et al., 1989; Freiburghaus et al., 1999). With the observation of the well-studied kilonova AT2017gfo (Tanvir et al., 2017) and the associated gravitational wave GW170817 and gamma ray burst GRB170817A (Abbott et al., 2017), recent research has been done on estimating the properties of the neutron star ejecta (Kasen et al., 2017; Bulla, 2023). This has led to developments in understanding the properties and equation of state of neutron stars (Radice et al., 2018; Mathias et al., 2024), additional measurements of the Hubble constant (Bulla et al., 2022), and investigating the contributions of these mergers to heavy element production via the r-process (Siegel, 2019).
@@
Over the past few years, developments in machine learning (ML) techniques have explored various avenues in reducing the computational cost and time for studying light curves. While ML-based improvements have been made in the vitesse of light curve generation using variational autoencoders (Saha et al., 2024) and neural networks (Peng et al., 2024), likelihood-free parameter estimation of this data remains an untapped field. A prohibitive step of sampling methods is the likelihood calculation, $p(d|\theta)$. To avoid this calculation, likelihood-free or simulation-based inference (LFI/SBI) offers a potential alternative method (see Cranmer et al. (2020) for a review).
In this paper, we describe our framework for faster parameter inference using LFI with comparable accuracy and precision to nested sampling (Skilling, 2006). We simulate light curves using the Ka2017 model grid and ZTF noise realizations. An embedding neural network marginalizes the distance to the source and the exact peak time over a data segment. This allows us to focus on the intrinsic parameters of the light curve such as the ejecta mass ($M_{\rm{ej}}$), velocity ($V_{\rm{ej}}$), and lanthanide fraction ($X_{\rm{lan}}$), that govern the light curve’s appearance. Further discussion of our choice of priors and simulation procedure are described in Section 2. We achieve marginalization by creating alternate, or shifted, views of each light curve with the same intrinsic parameters, while varying the time and luminosity distance. These views are jointly embedded by minimizing the variance, invariance, and covariance of the two data views. Conditioned on this data summary, a normalizing flow is used to infer these kilonova parameters on a logarithmic scale: $\log_{10}(M_{\rm{ej}}/M_{\odot})$, $\log_{10}(V_{\rm{ej}}/c)$, and $\log_{10}(X_{\rm{lan}})$. We compare our posterior distributions to nested sampling results produced with the PyMultiNest package (Feroz et al., 2009). Section 3 details the ML architecture and demonstrates its efficacy on simulated kilonova light curves, with the results provided in Section 4.
@@ 2 Data
We simulate light curve data using the Ka2017 radiative transfer method, which is available through the NMMA package (Kasen et al., 2017; Pang et al., 2024). We focus on obtaining realistic light curves viewable from ZTF, thereby simulating data in all three of its bands: ztfg, ztfr, and ztfi.
The composition and appearance of a kilonova is controlled by the spectral energy distribution (SED). The Ka2017 model parametrizes this SED through the combined dynamical and disk mass ($M_{\rm{ej}}$), velocity ($V_{\rm{ej}}$), and lanthanide fraction ($X_{\rm{lan}}$) of the ejecta. Lanthanide fraction, $X_{\rm{lan}}$, can serve as a metric for heavy r-process matter, as lanthanides contribute anywhere from 1 to 10 % of the total ejecta mass (Kasen et al., 2017). The line transitions from lanthanides greatly increases the opacity, resulting in long diffusion times and longer light curves. The Ka2017 model accounts for two kilonova mass ejection methods: a brief stripping of the stars due to tidal forces ($10^{-3}M_{\odot}<M_{\rm{ej}}<10^{-2}M_{\odot}$), and a secondary ejection of mass from the accretion disk blown by neutrino-driven winds ($10^{-2}M_{\odot}<M_{\rm{ej}}<10^{-1}M_{\odot}$) (Martin et al., 2015). The initial shock occurs at high speeds on the order of a dynamical velocity $V_{\rm{dyn}}\sim 0.2c-0.3c$, while the disk mass disperses at rates of $V_{\rm{disk}}\sim 0.05c-0.1c$ (Kasen et al., 2017). These quantities also affect the peak of the light curve: larger ejecta masses produce brighter and longer-lasting light curves, and faster ejecta velocities produce brighter yet short-lived light curves. Ka2017 combines these quantities as the ejecta parameters $M_{\rm{ej}}$ and $V_{\rm{ej}}$, which we use to parametrize our model. In this study, we neglect the effects of host galaxy and Milky Way extinction effects on the light curve.
@@
The investigation of kilonovae can be divided into two problems: search and inference. Here, we assume that ZTF is capable of discovering optical counterparts of gravitational waves to distances of up to 200 Mpc, and we simulate our light curves accordingly (Andreoni et al., 2021). To ensure that light curves are visible by ZTF, we selected broad parameter ranges that were capable of producing detections in addition to the considerations provided by Kasen et al. (2017). These priors are detailed in Table 1. Prior to training the model, all five parameters are normalized between 0 and 1. We choose time and luminosity distance, $(t,d_{L})$, as our nuisance parameters that we wish to marginalize. Our time prior is selected to create a conservative variability for where the light curve peaks. Our distance prior allows our model to be sensitive to the edge of EM detectablility with ZTF and the expected two detector LIGO O4 BNS inspiral range of 25-130 Mpc (Abbott et al., 2018). While GW170817 was detected at 40 Mpc, optical surveys are expected to be sensitive to kilonovae at a distance of up to 200 Mpc with a limiting magnitude of 22 mag (Sagués Carracedo et al., 2021).
@@
Our data is simulated at a regular cadence of one detection every six hours over a 20 day period of time. While this is an optimistic detection scheme, we discuss potential relaxation schemes in Section 4. We sample uniformly from our intrinsic parameters $\log_{10}(M_{\rm{ej}}/M_{\odot})$, $\log_{10}(V_{\rm{ej}}/c)$, and $\log_{10}(X_{\rm{lan}})$, and our extrinsic parameters $t$ and $d_{L}$. In the case where the light curve is dim or short-lived, non-detections are encoded as the magnitude detection limit of 22.0 in all bands. This is a typical conservative restriction given that ZTF is limited to 20.8, 20.6, and 19.9 mag [21.6, 21.4, and 21.9 mag] for the g, r, and i bands for a 30s [300s] exposure time (Bellm et al., 2019; Almualla et al., 2021). Since the broad priors produce both extremely short and long light curves, we pad the data with our non-detection limit in order to maintain a standard length of data for our model. For training purposes, a minimum amount of 8 cumulative detections across the three bands is set as a requirement.
Additionally, each light curve is simulated with unique instrument noise to ensure that all views of the same intrinsic parameters are distinct. These noise realizations come from the three noise distributions of the ztfg, ztfr, and ztfi bands. The noise instances are sorted into 12., 18., 20., 21., and 23 magnitude bins to obtain an accurate noise sampling for each data point in the light curve. The final processing step is a normalization performed across our training and test data. Neglecting the non-detections, we compute the mean and standard deviation of our training sample and apply the following procedure to normalize the photometric data:
$z_{i}=\frac{x_{i}-\bar{x}}{\sigma},$ (1)
where $x_{i}$ represents our original data and $z_{i}$ the new normalized photometry.
@@
To train the network, we create a set of “fixed” data, $d$, which share the same initial arrival time $t=0$ and a luminosity distance of $d_{L}=50$ Mpc. For each fixed light curve, an alternative “shifted” light curve $d^{\prime}$ is generated by modifying $t$ and $d_{L}$ according to the priors in Table 1. A single fixed light curve will have its shifted counterpart that shares the precise values for $\log_{10}(M_{\rm{ej}}/M_{\odot})$, $\log_{10}(V_{\rm{ej}}/c)$, and $\log_{10}(X_{\rm{lan}})$, resulting in the same physical phenomena but a different observed light curve. In this way, we create two views of the data generated by the same intrinsic parameters with distinct noise instances. We utilize both views to train our data summary, and only the shifted view to train our normalizing flow.
@@ 3 Method
In this section, we present our ML-based framework for parameter estimation using LFI. We introduce the concept of an embedding network that permits the marginalization of certain parameters and reduces the size of the following inference network in Section 3.1. The embedding also serves to reduce the high-dimensional light curves, each consisting of 121 points multiplied by three bands, to a 7 dimensional representation. The pre-trained embedding is supplied to a normalizing flow, which performs the parameter inference as described in Section 3.2.
@@ 3.1 Data Summary
In order to perform LFI using normalizing flows, we sample parameters $\boldsymbol{\theta}_{i}\sim p(\boldsymbol{\theta})$, and simulate light curves $\boldsymbol{d}_{i}\sim p(\boldsymbol{d}\mid\boldsymbol{\theta}_{i})$. The resulting pairs $\{\boldsymbol{\theta}_{i},\boldsymbol{d}_{i}\}$ are used to learn an approximation of the posterior distribution, $q_{\phi}(\boldsymbol{\theta}\mid\boldsymbol{d})$, by maximizing the log-likelihood,
$\log\mathcal{L}(\phi)=\sum_{i}\log q_{\phi}(\boldsymbol{\theta}_{i}\mid\boldsymbol{d}_{i}).$ (2)
where $q_{(\phi)}$ is the model’s estimated posterior.
A key challenge in implementing LFI on our dataset is the high dimensionality of our light curves. To combat this, we train an effective data summary that serves a dual purpose: reducing data dimensionality and marginalizing nuisance parameters: time and distance. In this case, we are primarily interested in the intrinsic parameters, like $M_{\rm{ej}}$, $V_{\rm{ej}}$, or $X_{\rm{lan}}$, though the data depends on other extrinsic parameters. Kilonova signals can be seen as unique if they arrive at different times or occur at different distances, even when the intrinsic parameters of the ejecta are the same. The recorded peak time of a kilonova depends not only on its arrival time and luminosity distance, but also on the underlying model, providing yet another reason to marginalize over it. An LFI network unaware of these translations requires the extrinsic parameters of time and distance to be inferred alongside the other physical parameters or marginalized over, leading to a larger, more expensive network. To mitigate the size of the network, we can marginalize the nuisance parameters by creating an appropriate data summary that preserves the information of only the intrinsic parameters and learns the symmetric patterns of $t$ and $d_{L}$ through different views of the data. We utilize self-supervised learning (SSL) to accomplish this goal.
This technique has been used previously in Chatterjee et al. (2023) for simpler, lower-dimensional signal morphologies, demonstrating that the creation of an embedding space aids in faster parameter inference. The application of self-supervised learning to create a summary extends to compressing cosmological data and encoding symmetries of partial differential equations as seen in Akhmetzhanova et al. (2024); Mialon et al. (2023). Additionally, self-supervised learning has been used to simplify the complexity of gravitational waves as an initial denoising step (Lin & Kong, 2024). Rodrigues & Gramfort (2020) demonstrates a different approach (YuleNet) for data summarization for LFI with timeseries data. As a result, we also implement a summarization step for our high-dimensional light curve data.
@@
Our full network architecture is explained in Figure 2’s flowchart. After the representation has been extracted, we pass $\gamma$ and $\gamma^{\prime}$ to our 150-dimensional expander layer, $h$, such that the outputs $x,x^{\prime}$ are produced:
@@
Here, $x$ and $x^{\prime}$ represent the data in an embedding space, which we denote as the data summary. Our performance hinges on a sizable $h(\gamma)$. Implementing a large, fully-connected linear layer allows information distributed across several nodes to flow between one another. This enables the network to condense information across the entire light curve.
@@
To correlate $x$ and $x^{\prime}$, we utilize the VICReg loss function (Bardes et al., 2021), which is calculated as a weighted sum on batches $X=[x_{1},...,x_{n}]$ and $X^{\prime}=[x_{1}^{\prime},...,x_{n}^{\prime}]$:
@@
$\begin{split}\mathcal{L}_{VICReg}(X,X^{\prime})&=\lambda_{1}MSE(X,X^{\prime})+\lambda_{2}[\sqrt{Var(X)+\epsilon}\\ &+\sqrt{Var(X^{\prime})+\epsilon}]+\lambda_{3}[C(X)+C(X^{\prime})],\end{split}$ (5)
where $C(X)$ is the covariance matrix, defined by:
$C(X)=\frac{1}{n-1}\sum^{n}_{i=1}(x_{i}-\bar{x})(x_{i}-\bar{x})^{T}$ (6)
where
$\bar{x}=\frac{1}{n}\sum^{n}_{i=1}x_{i}.$ (7)
@@
We choose equal weights $\lambda_{1}=\lambda_{2}=\lambda_{3}=1$ for this study. We minimize the VICReg similarity loss between the embeddings $x$ and $x^{\prime}$ such that the regularized variance and covariance of the batches are individually reduced, while the invariance between batches is reduced through mean-squared error. 10 minutes of processing time is required to train the network from scratch for 50 epochs on one NVidia V100 GPU.
@@
To properly encode the effects of changing $t$ and $d_{L}$, we must create two views of the data. We repetitively create 50 light curves using one identical combination of our physical parameters ($\log_{10}(M_{\rm{ej}}/M_{\odot})$, $\log_{10}(V_{\rm{ej}}/c)$, and $\log_{10}(X_{\rm{lan}})$) with unique noise instances. This is done iteratively with novel combinations of the physical parameters to produce a dataset. Each of light curves in this fixed set are positioned at a $d_{L}$ of 50 Mpc, and start at the same time. Our second data view is created by adjusting $t$ and $d_{L}$ of each light curve of the fixed set according to the priors listed in Table 1. Our training dataset contains 8,729 unique combinations of $M_{\rm{ej}}$, $V_{\rm{ej}}$, and $X_{\rm{lan}}$, resulting in 872,900 total light curves. The data representation, $\gamma$, is trained jointly using these sets of fixed and shifted light curves.
Figure 3: In this figure, we see four combinations of $\log_{10}(M_{\rm{ej}}/M_{\odot})$, $\log_{10}(V_{\rm{ej}}/c)$, and $\log_{10}(X_{\rm{lan}})$ plotted in orange, green, red, and purple. Each set contains 1000 light curves that have been shifted in time and distance. These plots represent a data summary space. The left plot displays a pre-trained embedding, and the right plot displays the embedding after it has continued training alongside the normalizing flow. By passing an informed embedding where parameters have been clustered and allowing it to further be trained, we find that the embedding space develops a more complex mapping to further separate different light curves while maintaining clustering of distinct $\log_{10}(M_{\rm{ej}}/M_{\odot})$, $\log_{10}(V_{\rm{ej}}/c)$, and $\log_{10}(X_{\rm{lan}})$.
$\log_{10}(M_{\text{ej}})$ $\log_{10}(V_{\text{ej}})$ $\log_{10}(X_{\text{lan}})$
Set 1 $-1.00>M>-1.25$ $-0.53>V>-0.75$ $-3.00>X>-4.50$
Set 2 $-1.25>M>-1.50$ $-0.75>V>-1.00$ $-4.50>X>-6.00$
Set 3 $-1.50>M>-1.75$ $-1.00>V>-1.25$ $-6.00>X>-7.50$
Set 4 $-1.75>M>-1.90$ $-1.25>V>-1.52$ $-7.50>X>-9.00$
Table 2: Parameter ranges for the four ejecta clusters shown in the plot above. Each set corresponds to a distinct injection of $\log_{10}(M_{\rm{ej}}/M_{\odot})$, $\log_{10}(V_{\rm{ej}}/c)$, and $\log_{10}(X_{\rm{lan}})$, with rows color-coded to match their respective cluster colors in the plot.
@@
Figure 3 displays the representation $\gamma^{\prime}$ for four combinations of physical injection parameters to showcase various areas of the prior selection. A visual comparison of the left and right plots highlights the importance of allowing a portion of the embedding weights to remain trainable alongside the normalizing flow. We observe that after training, light curves of different $\log_{10}(M_{\rm{ej}}/M_{\odot})$, $\log_{10}(V_{\rm{ej}}/c)$, and $\log_{10}(X_{\rm{lan}})$are clustered even with different values of $t$ and $d_{L}$, indicating that the two invariant parameters have been successfully marginalized.